OpenStudy (anonymous):

The width of a rectangle is 2cm less than its length. the perimeter of the rectangle is 16 cm. What is the length of the rectangle? show work please

OpenStudy (anonymous):

If the perimeter is 16, then 16 must be split up into 4 values. Where two are x and the other two are x-2. If you have a quick look at 16, try to split it up. E.g. 6 = 3+3 and 10 = 5 + 5 Then the lenght is 5 and the width is 3

OpenStudy (anonymous):

You could also use algebra, however, this is just as simple.

OpenStudy (anonymous):

p = 2l + 2w 16 = 2l + 2(l-2) 16/2 = l + (2(l-2))/2 8 = l + l - 2 10 = l + l 10 = 2l 10/2 = l 5 = l w = l-2 w = 3, l=5

OpenStudy (anonymous):

There you go.

OpenStudy (anonymous):

thanks